Session: Oct 2005 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Tao An at Wed Nov 2 14:28:36 2005) Success :

Urumqi (Zhang hua at Wed Nov 2 17:37:58 2005) Success :

Onsala (Michael Lindqvist at Wed Nov 2 18:33:10 2005) Success :
Only one (1) tsys in the schedule. The experiment consists of a single scan ==> Problem with disk change, see mail from Alastair.

Torun (Kaz Borkowski at Thu Nov 3 06:40:51 2005) Minor failures :
Off source from 12:17 to 12:31 due to an electricity outage. No other known problems (the single big file has luckily fit on current diskpack).

Medicina (Andrea Orlati at Thu Nov 3 11:02:00 2005) Success :
Experiment contains one huge scan (1.6 TB). We were forced to do a bank switch manually. No other problems were spotted.

Westerbork (Tony Foley at Fri Nov 4 10:29:58 2005) Success :
Wb used in phased array mode. No known problems

Effelsberg (A. Kraus at Wed Nov 9 16:36:17 2005) Minor failures : RECO HIGH
Same problem as Mc, On, JB with the one single scan. Problems when bank switch was due: data lost between UT0432-0641. The only scheduled T_sys measurement in the beginning was corrupt as the telescope was partly shadowed by surrounding hills. Added a second measurement at the end.

Jodrell_Bank (Alastair Gunn at Thu Nov 10 12:28:25 2005) Minor failures : WIND SCHED
The Lovell Telescope was used. Telescope parked due to high winds between 02:52 and 04:41 UT. Telescope was not making source changes in time during zenith transits. No other known problems. Managed to find a disk pack large enough for the single 1.6 TB scan!